2. (10 Points) Use source transformation to find V. 10 w + V- + 8 V 222 ЗА
Answer #1

(2 m + V- 2-12 BA + BV my 22 +372 3A = 6V 12 2-2 we + + V- I 2 6V Apply KUL -8 +1xI) + QxI) + 6 = 31 = 2 I = A WlN IX I VE No
Answer #2

(2 m + V- 2-12 BA + BV my 22 +372 3A = 6V 12 2-2 we + + V- I 2 6V Apply KUL -8 +1xI) + QxI) + 6 = 31 = 2 I = A WlN IX I VE No

answered by: ANURANJAN SARSAM